# QuantFactor REINFORCE: Mining Steady Formulaic Alpha Factors with Variance-bounded REINFORCE

### v2

@@ III-C MDPs for Mining Formulaic Alpha Factors
Expressions that are syntactically correct might still fail to evaluate due to the restrictions imposed by certain operators. For example, the logarithm operator token is not applicable to negative values. Such invalidity can not be directly detected. Therefore, these expressions are assigned a reward of $-1$ (the minimum value of IC) to discourage the policy from generating these expressions.
@@ IV-A The Proposed Algorithm
The detailed pipeline of the QFR algorithm is shown in Fig. 3. The policy model generates one new token (action) at each step, and the sequence of generated tokens (states) can be uniquely converted into an expression based on RPN. The policy model uses random sampling and greedy sampling to complete two trajectories and generate two expressions, i.e., two factors. The combination model is used to maintain a weighted combination of principal factors and, at the same time, to evaluate these two factors. Specifically, the combined factor values are used to calculate $r\left(\mathbf{a}_{1:T}\right)$ and $r\left(\mathbf{\bar{a}}_{1:T}\right)$ along with the real asset data features $\mathcal{X}$. These two rewards, calculated by the proposed shaped reward function (detailed in Section IV-B), form the estimation of (8). The algorithmic steps are given in Algorithm 1.
@@ IV-B Time-varying Reward Shaping
The reward function in [19] focuses on the absolute excess returns of factors, while ignoring the risk-adjusted characteristics of the returns, i.e., how the factors are resistant to the volatility of the market. In order to better balance returns and risks, our work not only evaluate the predictive accuracy of factors, but considers the stability of their predictive signals, thus providing a more comprehensive assessment of factor performance.
@@
Equation (11) is the proposed shaped reward. $\mathbb{I}$ denotes the indicator function, which takes the value of $1$ if the condition inside the parentheses is satisfied, and 00 otherwise. $\alpha$ is the time delay, set to $9\times 10^{4}$ time steps, $\eta$ is the slope of the change in the IR test value, set to $2.65\times 10^{-6}$, $\delta$ is the maximum IR test value, set to 0.3, and $\lambda$ is used to balance the main reward with the shaped reward, set to 0.02. The varying of the clip function over training time is illustrated in Fig. 4.
Using the IR test as reward shaping can force the QFR algorithm focusing more on the long-term stability and consistency of the factors during optimization, rather than just the short-term profit peaks. This reward shaping method guides the policy network to find factors that maintain stable predictive capability across different market conditions, reducing the risk of overfitting the training data. It helps to reduce the model’s over-reliance on specific market conditions, thereby enhancing the generalization ability of factors across different market environments.
@@ IV-C The Theoretical Analysis
The proof is thus completed. ∎
@@
will ensure $\operatorname{Var}[\tilde{g}(\mathbf{\theta})]<\operatorname{Var}[\hat{g}(\mathbf{\theta})]$. Notably, (24) reveals that if $r_{1}<2r_{2}$, any possible $p\in(0,1)$ guarantees a lower variance using QFR. The result can be generalized under a larger action space, and shows an advantage of QFR when the optimal action is not fully dominated. ∎
@@ V Numerical Results
In this section, we numerically evaluate QFR, comparing it with both state-of-the-art RL algorithms and other factor mining methods. Our experimental study is comprised of six stock datasets (detailed in Section V-A), and evaluates the efficiency of various RL algorithms in solving MDPs for mining formulaic alpha factors in Section V-B, after which we study the hyper-parameters in reward shaping in Section V-C, and considers five factor mining methods (discussed in Section V-D and V-E). Finally, the ablation study confirms the importance of the two improvements in Section V-F.
@@ V-A Environment Settings
The raw data sourced from the Chinese A-shares market, as well as the US stock market, specifically focusing on the constituent stocks of the China Securities Index 300 (CSI300), the China Securities Index 500 (CSI500), the China Securities Index 1000 (CSI1000), the S&P 500 Index (SPX), the Dow Jones Industrial Average (DJI), and the NASDAQ 100 Index (NDX) are utilized to model the MDPs for mining formulaic alphas in our experiment. To ensure reproducibility, we have only identified six primary features to generate the formulaic alphas, which include opening price (open), closing price (close), highest price (high), lowest price (low), trading volume (volume), and volume-weighted average price (vwap). Our objective is to generate formulaic alphas that exhibit a high IC with respect to the ground-truth 5-day asset returns. The dataset is split into three subsets: a training set spanning from 2016/01/01 to 2020/01/01, a validation set spanning from 2020/01/01 to 2021/01/01, and a test set spanning from 2021/01/01 to 2024/01/01. All price and volume data have been forward-dividend-adjusted respected to the adjustment factors on 2023/01/15.
To evaluate how well our framework performs against traditional formulaic alpha generation approaches, tree models, heuristic algorithms, end-to-end deep model algorithms, and interpretable reinforcement learning are adopted as baseline algorithms. We follow the open-source implementations of AlphaGen[19], gplearn[58] Stable Baseline 3[59] and Qlib[41] to produce the results.
@@
    • AlphaGen[19]: A natural solution to the MDPs for mining formulaic alphas, utilizing reinforcement learning for the first time in finding such interpretable alpha factors.
In addition, both the QFR and AlphaGen are used for solving the MDPs defined in Section III-C. In order to demonstrate the state-of-the-art performance of QFR in this reinforcement learning task, some respected reinforcement learning baseline algorithms, including TRPO[60], PPO[21] and A3C[61] are also utilized in the experiment.
Figure 5: Performance of our method and state-of-the-art reinforcement learning methods for mining formulaic alpha factors in the constituent stocks of CSI300, CSI500, CSI1000, SPX, DJI, and NDX. All the curves are averaged over 5 different random seeds, and half of the standard deviation is shown as a shaded region.
To demonstrate the effect caused by stochasticity in the training process, each experimental combination with an indeterministic training process is evaluated with 5 different random seeds. The hyperparameters of MLP, XGBoost and LightGBM are set according to the benchmarks given by Qlib. The hyperparameters of GP are set according to the gplearn framework. The actor network and critical network of PPO, A3C and TRPO share a base LSTM feature extractor, which has a 2-layer structure with a hidden layer dimension of 128. The drop out rate is set to 0.1. The separate value and policy heads are MLPs with two hidden layers of 64 dimensions. PPO clipping range $\epsilon$ is set to 0.2. TRPO trust region constraint upper bounds $\delta$ is set to $1.1\cdot 10^{-5}$. All the operator tokens used in our experiment are shown in Table IV. The simulation is performed by a single machine with an Intel Core i9-13900KCPU and two NVIDIA GeForce RTX 4090 GPUs.
TABLE IV: All the Operator Tokens Used in the Experiment
@@ V-B Comparisons with Other RLs
When solving the MDPs defined in Section III-C, QFR demonstrates superior performance compared to PPO that is adopted in AlphaGen[19]. We present experimental results on the constituent stocks of six indices in China and the United States using typical state-of-the-art RL algorithms, including TRPO, PPO, and A3C, as shown in Fig 5. Due to the fact that QFR utilizes reward shaping, while the baseline algorithm does not, reward is not used as a metric when comparing the learning curves of various algorithms. Instead, Rank Information Coefficient (Rank IC) is used to substitute for reward in evaluating the performance of the policy network during training. The Rank IC assesses the association between the ordinal positions of alpha values and the sequences of stock return rankings. Rank IC is just the IC of ranked data, defined as $RankIC\left(\mathbf{z}^{\prime}_{t},\mathbf{y}_{t}\right)=IC\left(r(\mathbf{z}^{\prime}_{t}),r(\mathbf{y}_{t}\right))$, where $r(\cdot)$ is the ranking operator. This metric is the higher the better. It can be observed that our method can outperform the other RL algorithms in all six indices, especially at the end of the training process. Compared to the best-performing PPO algorithm adopted by AlphaGen, our algorithm improves by 3.83% on the metric, which further validates the effectiveness of discarding elements related to the critic network, introducing a subtractive baseline value, as well as the reasonable reward shaping.
V-C The Studies of the Time-varying Reward Shaping
@@
Figure 6: Performance of QFR with different slope $\eta$, time delay $\alpha$, and maximum IR test value $\delta$, given the same number of iterations.
We study the impact of variations in the slope $\eta$, time delay $\alpha$, and maximum IR test value $\delta$ in reward shaping on the learning process. As shown in Fig. 6, we find that a larger $\eta$ tends to achieve high performance. However, when $\eta$ is too large, the intensity of IR test becomes excessive, causing the policy network to continuously receive discounted rewards after $\alpha$ steps, which harms the efficiency of QFR. Therefore, an appropriate slope of $2.65\times 10^{-6}$ can better trade off between performance and efficiency. The time delay $\alpha$ becomes smaller as the alpha factor quality improves, but an $\alpha$ less than $7\times 10^{4}$ will have a negative impact on the alpha factor quality. We speculate that after the alpha factor quality becomes acceptable, it is appropriate to further optimize the factor quality using IR test mechanism, whereas introducing IR test too early in the training process may not have this effect. An interesting behavior to note is that the training process is highly robust to the maximum IR test value $\delta$. We settled on a value of $0.3$ for this parameter.
@@ V-D Factors Evaluation
We applied the baseline algorithms and our QFR algorithm to the constituent stocks of the CSI300 and CSI500 indices, and evaluated their performance based on two metrics: IC and Rank IC. The IC is delineated in (2). Both of the metrics are the higher the better. Note that all the algorithm is only optimized against the IC metric. The results are shown in Table V. The performance of MLP, XGBoost, and LightGBM is inferior due to the usage of the open-source set of alphas. On the other hand, AlphaGen tends to converge to local optima, leading to overfitting on the training data. Although GP mitigates this issue by preserving a diverse population, it still struggles to generate alphas that exhibit synergy in combination. The QFR algorithm excels in solving MDPs for mining formulaic alpha factors compared to AlphaGen. We attribute the superior performance of QFR to the acceleration of the convergence process by discarding elements related to the critic network, as well as the reasonable reward shaping.
Figure 7: Backtest results on CSI 300. The lines track the profit and loss of simulated trading strategies utilizing the various alpha factors minned from different algorithms. And the value of the y axis represents the cumulative return.
Figure 8: Performance of the mined factors on the CSI300 under different market volatility conditions. Three significant event-driven volatility anomaly periods were selected to demonstrate the ability of each algorithm to withstand volatility. The three periods are characterized by high market volatility driven by the Establishment of the Beijing Stock Exchange, a rapid decline in market volatility driven by the Hiking of interest rates from the Fed, and low market volatility lingering around driven by Approval-based IPOs.
V-E Investment Simulation
To further demonstrate the effectiveness of our algorithm in more realistic investing settings, we conduct investment simulation using the factors extracted by various algorithms. Specifically, these factors are applied to an index enhancement strategy, where stocks are first sorted according to their factor values, then the top-50 stocks are selected and a rebalancing operation is performed (selling those in the current portfolio that are not in the top-k stocks, and buying those in the top-k stocks that are not currently in the portfolio). We conducted backtest in the testing period (2021/01/01 to 2024/01/01) on the CSI300 index. The performance of the factors in the backtest was evaluated using cumulative returns (profit and loss), with the final value of this metric being the higher, the better. The result of the backtest is shown in Fig. 7. Although our algorithm does not consistently have the best performance throughout the entire backtesting process, it still performs well at the end and achieves the highest cumulative returns compared to the other baseline algorithms.
TABLE V: Performance of Mined Factors on CSI300 and CSI500
Method CSI300 CSI500

### v3

@@ III-C MDPs for Mining Formulaic Alpha Factors
Notably, this MDP only models the formulaic alpha factors generation process, and the environment here specifically refers to the QFR’s RL environment, rather than the financial market’s random behavior. Its deterministic feature is manually designed and holds true regardless of the specific characteristics of the financial market.
@@ IV-A The Proposed Algorithm
The detailed pipeline of QFR is shown in Fig. 3. The policy model generates one new token (action) at each step, and the sequence of generated tokens (states) can be uniquely converted into an expression based on RPN. The policy model uses random sampling and greedy sampling to complete two trajectories and generate two expressions, i.e., two factors. The combination model is used to maintain a weighted combination of principal factors and, at the same time, to evaluate these two factors. Specifically, the combined factor values are used to calculate $r\left(\mathbf{a}_{1:T}\right)$ and $r\left(\mathbf{\bar{a}}_{1:T}\right)$ along with the real asset data features $\mathcal{X}$. These two rewards, calculated by the proposed shaped reward function (detailed in Section IV-B), form the estimation of (8). The algorithmic steps are given in Algorithm 1.
@@ IV-B Time-varying Reward Shaping
The reward function in [14] focuses on the absolute excess returns of factors, while ignoring the risk-adjusted characteristics of the returns, i.e., how the factors are resistant to market volatility. In order to better balance returns and risks, QFR not only evaluate the predictive accuracy of factors, but considers the stability of their predictive signals, thus providing a more comprehensive assessment of factor performance.
@@
Equation (11) is the proposed shaped reward. $\mathbb{I}$ denotes the indicator function, which takes the value of 1 if the condition inside the parentheses is satisfied, and 0 otherwise. The clip function is defined as $\text{clip}(x,a,b)=\min\left(\max(x,a),b\right)$, which is utilized to constrain the range of values. $\alpha$ is the time delay, $\eta$ is the slope of the change, $\delta$ is the maximum IR test value, and $\lambda$ is the shaping coefficient. The varying of the clip function over training time is illustrated in Fig. 4.
Using the IR test as reward shaping can force QFR focusing more on the long-term stability and consistency of the factors during optimization, rather than just the short-term profit peaks.
@@ IV-C The Theoretical Analysis
The proof is thus completed. Notably, the noise in real asset data features $\mathcal{X}$ potentially increases $r_{\text{max}}$. This could result in an increase in the upper bound of $8r_{\text{max}}^{2}T^{2}/N$, but the proposition remains correct. ∎
@@
will ensure $\operatorname{Var}[\tilde{g}(\mathbf{\theta})]<\operatorname{Var}[\hat{g}(\mathbf{\theta})]$. Notably, (24) reveals that if $r_{1}<2r_{2}$, any possible $p\in(0,1)$ guarantees a lower variance using QFR. This result extends naturally to larger action spaces: variance reduction occurs when $r\left(\mathbf{a}_{1:T}\right)>r\left(\overline{\mathbf{a}}_{1:T}\right)/2$. Given that industry benchmarks for high-quality factors rarely exceed ICs of $0.1$[52], this condition holds even in the worst case once $r\left(\mathbf{a}_{1:T}\right)>0.05$. ∎
@@ V Numerical Results
In this section, we numerically evaluate QFR, comparing it with both state-of-the-art RL algorithms and other factor mining methods. Our experimental study is comprised of six stock datasets (detailed in Section V-A), and evaluates the efficiency of various RL algorithms in solving MDPs for mining formulaic alpha factors in Section V-B, after which we study the hyper-parameters of reward shaping in Section V-C, and considers five factor mining methods (discussed in Section V-D and V-E). Finally, the ablation study confirms the importance of the two improvements in Section V-F.
@@ V-A Environment Settings
The raw data sourced from the Chinese A-shares market, as well as the US stock market, specifically focusing on the constituent stocks of the China Securities Index 300 (CSI300, the index composed of the 300 most liquid and largest A-share stocks listed on the Shanghai and Shenzhen Stock Exchanges), the China Securities Index 500 (CSI500, the index representing 500 A-share stocks with mid-cap market values), the China Securities Index 1000 (CSI1000, an index including 1000 smaller-cap A-share stocks), the S&P 500 Index (SPX, the Dow Jones Industrial Average, which tracks 30 major US blue-chip companies representing key sectors), the Dow Jones Industrial Average (DJI), and the NASDAQ 100 Index (NDX, the NASDAQ-100, consisting of 100 of the largest non-financial companies listed on the Nasdaq Stock Market) are utilized to model the MDPs for mining formulaic alphas in our experiment. Given many features including macroeconomic features, company’s fundamental features, and stock price-volume features are not always publicly available, we have only identified six publicly available primary stock price-volume features for reproducibility to generate the formulaic alphas. They include the opening price (open), closing price (close), highest price (high), lowest price (low), trading volume (volume), and volume-weighted average price (vwap). Our objective is to generate formulaic alphas that exhibit a high IC with respect to the ground-truth 5-day asset returns. The dataset is split into three subsets: a training set spanning from 2016/01/01 to 2020/01/01, a validation set spanning from 2020/01/01 to 2021/01/01, and a test set spanning from 2021/01/01 to 2024/01/01. All price and volume data have been forward-dividend-adjusted respected to the adjustment factors on 2023/01/15.
To evaluate how well our framework performs against existing methods, tree models, heuristic algorithms, end-to-end deep model algorithms, and interpretable RL are adopted as baseline algorithms. We follow the open-source implementations of AlphaGen[14], gplearn[53] Stable Baseline 3[54] and Qlib[39] to produce the results.
@@
    • AlphaGen[14]: A natural solution to the MDPs for mining formulaic alphas, utilizing RL for the first time in finding such interpretable alpha factors.
In addition, both QFR and AlphaGen are used for solving the MDPs defined in Section III-C. In order to demonstrate the performance of QFR in this RL task, some respected RL algorithms, including TRPO[56], PPO[16] and A3C[57] are also utilized in the experiment.
Figure 5: Training-phase correlation between mined factor values and the prices of the six index constituents for the learning time of all investigated RL algorithms. All the curves are averaged over 5 different random seeds, and half of the standard deviation is shown as a shaded region.
To demonstrate the effect caused by stochasticity, each experiment is conducted with 5 different random seeds. The hyperparameters of MLP, XGBoost and LightGBM are set according to the benchmarks given by Qlib. The hyperparameters of GP are set according to the gplearn framework. The actor network and critical network of PPO, A3C and TRPO share a base LSTM feature extractor, which has a 2-layer structure with a hidden layer dimension of 128. The drop out rate is set to 0.1. The separate value and policy heads are MLPs with two hidden layers of 64 dimensions. PPO clipping range $\epsilon$ is set to 0.2. TRPO trust region constraint upper bounds $\delta$ is set to $1.1\cdot 10^{-5}$. All the operator tokens used in our experiment are shown in Table III. The experiments is performed by a single machine with an Intel Core i9-13900K CPU and two NVIDIA GeForce RTX 4090 GPUs.
TABLE III: All the Operator Tokens Used in the Experiment
@@ V-B Comparisons with Other RLs
We present the experimental results of state-of-the-art RL algorithms (including TRPO, PPO, and A3C) for mining formulaic factors on six indices, as shown in Fig. 5. Due to the fact that QFR utilizes reward shaping, while the other algorithms do not, Rank Information Coefficient (Rank IC) rather than reward is used as a metric when comparing the learning curves. Rank IC is just the IC of ranked data, defined as $RankIC\left(\mathbf{z}^{\prime}_{t},\mathbf{y}_{t}\right)=IC\left(r(\mathbf{z}^{\prime}_{t}),r(\mathbf{y}_{t}\right))$, where $r(\cdot)$ is the ranking operator. This metric is the higher the better. The six indices are constructed from stocks traded on diverse exchanges, including the Shanghai Stock Exchange, the Shenzhen Stock Exchange, the Nasdaq Stock Market, and the New York Stock Exchange. In effect, these stocks are very different: the CSI300 constituents tend to be large, well-established companies; the CSI500 constituents mainly comprise mid-cap stocks; the CSI1000 constituents is dominated by small- or lower-cap firms; the NDX constituents feature technology-driven and innovative companies; the DJI constituents represent established industry leaders; and the SPX constituents cover a wide array of large-cap US corporations. These different constituent stocks may reflect varying levels of market stress in different periods. For instance, NDX constituents tend to exhibit higher market stress during tech downturns. According to Fig. 5, apart from the experiment on CSI500 constituents where QFR and PPO exhibit similar performance, QFR outperforms all other algorithms in convergence speed and stability and improves by 3.83% on the metric compared to the best-performing PPO algorithm adopted by AlphaGen on the remaining five index constituents. Specifically, QFR catches up to surpass initially superior of PPO/TRPO. The early advantage of PPO/TRPO is due to their controlled policy updates step size and therefore stability in update during cold-start phase. However, the restricted step sizes limit further improvements and make it harder to escape local optima. In contrast, without such restriction, QFR benefits from more accurate gradient estimations and ultimately surpasses PPO and TRPO after sufficient training. Considering that the training set from 2016/01/01 to 2020/01/01 spans four years of trading time and includes various stocks from six indices across four exchanges, encompassing various levels of market trading stress, this indicates that QFR also demonstrates superior performance under varying market trading stress. This enhanced performance can be attributed to QFR discarding elements associated with the critic network, utilizing a subtractive baseline, and implementing reasonable reward shaping.
@@
Figure 6: Sensitivity analysis of QFR to shaping coefficient ($\lambda$), time delay ($\alpha$), maximum IR test value ($\delta$), and slope ($\eta$).
V-C Sensitivity Analysis of the Time-varying Reward Shaping
We analyze the effects of reward shaping hyperparameters, including slope $\eta$, time delay $\alpha$, maximum IR test value $\delta$, and shaping coefficient $\lambda$. The dataset, by comprising the 300 largest and most liquid A-share stocks that collectively represent over 70% of the market capitalization, provides a representative foundation for evaluations. As shown in Fig. 6, increasing $\eta$ generally improves performance, but excessive values (e.g., $\eta>2.65\times 10^{-6}$ ) intensify IR tests, forcing the policy network to receive prolonged discounted rewards after $\alpha$ steps. We select $\eta=2.65\times 10^{-6}$ to balance IC and IR trade-offs. The time delay $\alpha$ decreases with enhanced factor quality, yet values below $7\times 10^{4}$ trigger premature IR test and destabilize the process. Setting $\alpha=9\times 10^{4}$ optimally activates IR test only after sufficient factor refinement. Notably, the maximum IR test value $\delta$ exhibits robustness, and we fix it at 0.3. For shaping coefficient $\lambda$, performance stays steady between $\lambda=0.02$ and $0.04$, but degrades significantly beyond $0.04$. For instance, $\lambda=0.06$ reduces IC by 2.27% and Rank IC by 1.12%, and $\lambda=0.1$ reduces IC by 3.36% and Rank IC by 5.73% compared to $\lambda=0.02$. While we explored adaptive $\lambda$ tuning, it ultimately underperformed the optimal static $\lambda=0.02$. We conservatively adopt $\lambda=0.02$ to avoid potential over‑amplification.
@@ V-D Factors Evaluation
We applied the baseline algorithms and QFR to the CSI300 and CSI500 indices, and evaluated their performance based on two metrics: IC and Rank IC. The IC is delineated in (2). Both of the metrics are the higher the better. Note that all the algorithm is only optimized against the IC metric. The results are shown in Table IV. The performance of MLP, XGBoost, and LightGBM is inferior due to the usage of the open-source set of alphas. On the other hand, AlphaGen tends to converge to local optima, leading to overfitting on the training data. Although GP mitigates this issue by preserving a diverse population, it still struggles to generate alphas that exhibit synergy in combination. QFR excels in solving MDPs for mining formulaic alpha factors compared to AlphaGen. We attribute the superior performance of QFR to the acceleration of the convergence process by discarding elements related to the critic network, as well as the reasonable reward shaping.
TABLE IV: Testing-Phase Correlation Between the Prices of CSI300/CSI500 Constituents and Factor Values Mined by All Investigated Factor Mining Algorithms
CSI300 CSI500
@@
TABLE V: Testing-Phase Backtesting Market Conditions and Risk Metrics of the Investment Strategy Driven by QFR and AlphaGen for Each Quarter on CSI300 Constituents
21Q1 21Q2 21Q3 21Q4 22Q1 22Q2 22Q3 22Q4 23Q1
Market Features
Highest
Daily
Volatility
Lowest
Daily
Volume
Highest
Daily
Volume
& Turnover
Normal
Market
Low
Bench
Return
Highest
Bench
Return
Lowest
Bench
Return
Lowest
Daily
Turnover
Lowest
Daily
Volatility
Daily Volatility 22.79% 21.06% 19.68% 17.94% 19.19% 21.71% 19.08% 20.68% 17.90%
(CIMV) (1.58%) (1.08%) (1.25%) (1.27%) (2.63%) (1.71%) (1.69%) (1.93%) (1.06%)
Daily Volume 6.96 6.70 9.13 7.85 8.11 8.31 7.31 6.93 6.94
(100 Million Lots) (0.89) (0.64) (1.48) (0.72) (1.03) (0.88) (1.04) (1.14) (0.90)
Daily Turnover 0.94 0.88 1.30 1.10 1.01 0.95 0.91 0.84 0.88
(Trillion Yuan) (0.16) (0.12) (0.16) (0.09) (0.13) (0.15) (0.16) (0.15) (0.12)
Market Conditions Bench Return -3.13% 1.21% -6.97% 1.52% -14.53% 6.21% -15.16% 1.75% 4.26%
QFR 7.55% -1.80% 21.06% 1.59% 0.48% 2.65% -6.08% 9.10% 9.74%
Cumulative Return AlphaGen 4.55% 9.48% 20.97% 0.21% -6.47% 8.30% -6.78% 2.09% 7.46%
QFR 1.912 -0.170 2.124 0.193 0.227 0.445 -0.689 1.201 1.405
Sharpe Ratio AlphaGen 1.116 2.700 3.432 0.106 -1.145 1.405 -1.560 0.553 2.418
QFR 5.88% 3.99% 4.60% 7.05% 10.33% 8.55% 7.80% 7.27% 4.38%
Maximum Drawdown AlphaGen 5.10% 3.72% 5.08% 7.14% 8.73% 8.16% 7.88% 6.56% 3.41%
QFR 183.71% 129.66% 605.62% 327.65% 439.15% 531.74% 114.90% 148.96% 79.94%
Strategy Risk Metrics Turnover AlphaGen 446.55% 508.80% 452.97% 321.84% 347.68% 385.03% 396.92% 445.46% 482.29%
V-E Investment Simulation
To further demonstrate QFR’s effectiveness in realistic investments, we conduct simulations using strategies driven by all investigated algorithm. The testing phase(2021/01/01–2024/01/01) employs CSI300 index, which tracks the 300 largest and most liquid A-share stocks. As China’s premier equity benchmark, it captures over 70% of the market’s total value, ensuring representative evaluations. The trading strategy is a top-50 strategy. On each trading day, the CSI300 constituents are first sorted according to the mined alpha factor values, then the top-50 stocks are selected for holding positions. If the top-50 stock list changes compared to the previous trading day, rebalancing is required (buying new entries and selling those removed), and we assume the difference between the expected and actual execution price is zero because daily trading using TWAP/VWAP strategies effectively neutralizes this kind of price slippage[58, 59, 60, 61, 62]. The performance of the investment strategy was evaluated using cumulative returns, with the final value of this metric being the higher, the better. The result of the backtest is shown in Fig. 7, Table V and Table VI.
Figure 7: Testing-phase backtesting results of the investment strategies driven by all investigated factor mining algorithms on CSI300 constituents. The lines track the cumulative returns of simulated investment strategies using the mined factors. The red background indicates that QFR dominates the compared methods with quarterly cumulative returns, while a green background indicates the contrary.
Figure 8: Testing-phase backtesting performance of the investment strategies driven by all investigated factor mining algorithms on CSI300 constituents under different market volatility conditions. Three significant event-driven volatility anomaly periods were selected to demonstrate the ability of each algorithm to withstand volatility.
Figure 9: Training-phase correlation between mined factor values and the prices of the six index constituents of for each learning episode of the two variants, together with QFR. All the curves in this ablation experiments are averaged over 5 different random seeds, and half of the standard deviation is shown as a shaded region.
It can be observed that QFR outperforms the baseline methods under most market conditions. Among 12 backtested quarters, the QFR-driven investment strategy failed to dominate in only 3 quarters. Even in periods of highest market daily volatility (21Q1), highest market daily volume and turnover (21Q3), lowest benchmark return (22Q3), lowest market daily turnover (22Q4), and lowest market daily volatility (23Q1), the QFR-driven investment strategy still outperforms the best baseline algorithm, AlphaGen, demonstrating its robust effectiveness under various market conditions and market stress. In the cumulative return comparisons across weekly, quarterly, and yearly time horizons, the QFR-driven investment strategy still outperforms all baseline algorithms, only slightly underperforming relative to AlphaGen in the monthly cumulative returns comparison, demonstrating its robust performance and consistent superiority over multiple time horizons.
TABLE VI: Testing-Phase Cumulative Returns of the Investment Strategies Driven by All Investigated Factor Mining Algorithms on CSI300 Constituents under Different Time Horizons
Cumulative Returns by Time Horizons (%)
Weekly Monthly Quarterly Yearly
MLP
0.124
0.387
1.733
7.158
[…]
0.128
0.423
1.702
6.790
[…]
0.104
0.333
1.402
5.538
[…]
0.127
0.636
2.115
8.292
[…]
0.238
1.114
3.620
14.515
QFR
0.247
0.999
4.277
18.673